OpenStudy (anonymous):

PLease help free medlas! Find the quotient 2/5 divided by -3/10

OpenStudy (anonymous):

what grade are you?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

reverse -3/10 and multiplied it to 2/5

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i dont like negitive numbers

OpenStudy (anonymous):

and it comes out as 1 and 1/3 but is it negitve of postive

OpenStudy (anonymous):

or*

OpenStudy (anonymous):

You can always convert 2/5 into x/10, and divide. 2/5 = 4/10. (4/10) / (-3/10) = -4/3

OpenStudy (anonymous):

Your answer should be -4/3

OpenStudy (anonymous):

Thanks Guys! i would give everyone that answered a medla but is wont let me so...

OpenStudy (anonymous):

Okay I would like one but if you cannot then that is fine! I am happy to help! :)